# ChainRank-DPO: Chain Rank Direct Preference Optimization for LLM Rankers

### v1

@@ 5 Limitations
While this study provides valuable insights, several limitations should be acknowledged. Firstly, the need to include long examples of CoT formatting (Step1: …) in our prompts leads to increased training and inference times.
Secondly, although LLaMA3 has a larger context window (8192 tokens) compared to Vicuna and Zephyr (4096 tokens), we are still limited by the maximum number of passages in our training data, which is set to 20. As a result, we are restricted to ranking within a window size of 20 and require nine passes to complete the reranking of 100 passages.
@@ Appendix A Baseline Selection
To illustrate the efficacy of ChainRank, we select several baseline models, including the existing retriever models, BM25 Robertson and Zaragoza (2009) and Contriever Izacard et al. (2021), zero-shot prompt-based LLMs, RankGPT4 and Gemma-7B Team et al. (2024), fine-tuned LLM-based models (RankVicuna, RankZephyr), as well as our base model LLaMA3-8B-Instruct. To show the effectiveness of our CoT reranking prompt, we train a model with a similar paradigm to RankVicuna but use our base model LLaMA3-8B-Instruct, denoted as RankLLaMA3.
Appendix B Window Size
Models Window/Stride Size 2/1 10/5 20/10 Gemma-7B 0.541 0.593 0.533 LLaMA3 0.549 0.660 0.641 RankZephyr 0.562 0.612 0.742 RankVicuna 0.551 0.655 0.668 RankLLaMA3 0.567 0.663 0.666 ChainRank-SFT 0.697 0.737 0.755 ChainRank-DPO 0.671 0.755 0.758
@@ Appendix B Window Size
Appendix C Explanation of CoT strategies
To evaluate the trade-off between performance and inference cost, we trained our model to generate ranking orders after every 1, 3, 5, and 7 steps. An additional instruction was introduced to guide the model in returning ranking orders at these specified intervals. For comparison, we also included the performance of RankLLaMA3 to highlight the effectiveness of our CoT strategies
Appendix D Examples of general LLM tasks
To illustrate the degradation in general text understanding and generation capabilities of RankVicuna and Rankzephyr, we provide two examples below. Due to the length of ChainRank’s response, we have omitted the middle portion of the answer.
D.1 Example 1
@@
ChainRank Aloha and Mahalo: Discovering the Aloha Spirit in Hawaii As I stepped off the plane in Honolulu, I couldn’t help but feel a sense of excitement and anticipation wash over me. I had dreamed of visiting Hawaii for years, and finally, I was here. The warm tropical air enveloped me, and the sweet scent of plumeria filled my senses as I made my way to my hotel. Little did I know, this trip would be a journey not only to breathtaking natural wonders but also to a rich cultural heritage that would leave me feeling grateful and inspired. Immersing in the Local Culture: (Content omitted for brevity.) Must-See Attractions: (Content omitted for brevity.) A Taste of Aloha: (Content omitted for brevity.) Conclusion: (Content omitted for brevity.) Practical Tips and Recommendations: Visit Pearl Harbor and USS Arizona Memorial in the morning to avoid the crowds. Try local specialties like poke, laulau, and shave ice. Take a surf lesson or snorkeling trip to explore the islands’ stunning coastlines. Visit the Bishop Museum to learn about the history and geology of Hawaii. Explore the islands’ many farmers’ markets for fresh produce and local products. Learn traditional Hawaiian crafts, like leis and ukulele strings, from local artisans. Mahalo (thank you) for joining me on this journey to the Aloha State. I hope my story has inspired you to plan your own trip to Hawaii and experience the warmth and beauty of this incredible destination.
D.2 Example 2
@@
ChainRank I’d be happy to help you write a resignation letter to your current employer. Here’s a sample template you can use as a guide: Dear [Recipient’s Name], I am writing to inform you of my decision to resign from my position as [Your Position] at [Company Name], effective [Date of Last Day of Work]. I want to express my deepest gratitude to you and the entire [Company Name] team for the opportunities I’ve had during my time here. (Content omitted for brevity.) Please don’t hesitate to reach out to me if you need anything after my departure. I can be reached at [Your Email Address] or [Your Phone Number]. Sincerely, [Your Signature] Some tips to keep in mind when writing your resignation letter: Be professional and respectful: Use a formal tone and language throughout the letter. Express gratitude: Let your employer know how much you appreciate the opportunities you’ve had and the experience you’ve gained. Offer to help with the transition: Show that you’re committed to making the transition as smooth as possible by offering to complete outstanding tasks and assist with the handover of your responsibilities. Keep it brief: Aim for a letter that is no more than one page in length. Proofread: Make sure to proofread your letter carefully to ensure it is free of errors and typos. Remember to customize the letter to fit your own circumstances and experiences. Good luck with your resignation, and I hope it leads to a positive and successful transition for you!

### v2

@@ 5 Limitations
While our model is based on LLaMA 3.1, which supports a 128k context window—significantly longer than RankVicuna (4k) and RankZephyr (8k)—we observe that using extremely long contexts by including all passages at once can negatively impact reranking performance. In some cases, this leads to worse results compared to a sliding window approach. This suggests that simply leveraging longer context lengths does not always yield better rankings, possibly due to attention dilution or ineffective utilization of distant context. In future work, we aim to explore strategies to better manage long-context inputs, such as adaptive passage selection or hierarchical attention mechanisms, to fully exploit the model’s context capacity without compromising ranking quality
@@ Appendix A Baseline Selection
To illustrate the efficacy of RaCT, we select several baseline models, including the unsupervised retriever model BM25 Robertson and Zaragoza (2009); zero-shot prompt-based LLMs: RankGPT4444gpt-4-0613, Gemma-7B Team et al. (2024), LLaMA3.1-8B-Instruct Grattafiori et al. (2024) and DeepSeek-R1-Distill-Llama-8B DeepSeek-AI et al. (2025); fine-tuned LLM-based models (RankVicuna, RankZephyr). To show the effectiveness of our CoT reranking prompt, we train a model with a similar paradigm to RankZephyr but use our base model LLaMA3.1-8B-Instruct, denoted as LLaMA3.1 + SFT.
Appendix B Evaluation Benchmarks (Expanded)
We evaluate ranking capabilities using TREC DL19 Craswell et al. (2020) and DL20 Craswell et al. (2021a) Tracks, derived from MS MARCO V1 Bajaj et al. (2016), with human-annotated relevance labels. DL19 and DL20 contain 43 and 54 queries, respectively, each paired with 100 BM25-retrieved candidate passages. For broader generalization analysis, we include additional datasets not seen during training: (1) BEIR Benchmarks Thakur et al. (2021), (2) TREC DL21 Craswell et al. (2021b) and DL22 Craswell et al. (2022) based on MS MARCO V2, and (3) BRIGHT: A Reasoning-Intensive Benchmark Su et al. (2025). Reranking uses a sliding window strategy Sun et al. (2023) with window size 20 and stride 10 to process long sequences. Rankings are computed independently for each segment and aggregated across the query.
To assess whether ranking fine-tuning affects broader model capabilities, we test on MMLU Hendrycks et al. (2020), which covers 57 diverse academic and professional subjects. We compare performance before and after ranking fine-tuning to ensure the model retains general language understanding and generation abilities.
Appendix C Sliding Window Strategy
In the sliding window strategy commonly used in reranking models (e.g., RankZephyr, RankGPT, and our method), passages appearing in multiple overlapping windows are ranked multiple times based on the local context of each window.
Since most reranking evaluations focus on nDCG@10, only the top 10 passages from the final ranking are considered. This means that passages appearing in the final window retain their latest assigned ranking, while earlier rankings from overlapping windows are not explicitly reconciled. This approach follows the standard strategy used in previous works and provides a computationally efficient way to handle long passage lists.
Appendix D Base Model Selection
Models DL19 DL20 DL21 DL22
LLaMA3-8B 0.752 0.710 0.693 0.502
LLaMA3.1-8B 0.753 0.710 0.698 0.504
LLaMA3.2-3B 0.718 0.697 0.690 0.517
LLaMA3.2-1B 0.713 0.682 0.673 0.465
Table 4: Performance Comparison of different base models on TREC Benchmarks (nDCG@10). All the models are trained with identical RaCT strategy. The bold values highlight the best performance across the respective datasets.
Appendix E Window Size
Models Window/Stride Size 2/1 10/5 20/10 Gemma-7B 0.541 0.593 0.533 LLaMA3.1 0.549 0.660 0.641 RankZephyr 0.562 0.612 0.742 RankVicuna 0.551 0.655 0.668 LLaMA3.1 + SFT 0.567 0.663 0.666 RaCT w/o RPO 0.697 0.737 0.753 RaCT 0.702 0.754 0.758
@@ Appendix E Window Size
Appendix F BEIR Evaluation
Datasets BM25 RankZephyr RaCT
NFC 0.322 0.331 0.392
FEVER 0.651 0.618 0.823
COVID 0.595 0.592 0.847
NEWS 0.395 0.396 0.471
SciFact 0.679 0.679 0.790
HotPotQA 0.633 0.669 0.762
NQ 0.306 0.375 0.600
Qoura 0.789 0.719 0.807
FIQA 0.236 0.230 0.435
CFEVER 0.165 0.160 0.237
Arguana 0.397 0.295 0.310
Touche 0.442 0.434 0.414
SCIDOCS 0.149 0.148 0.207
Signal 0.330 0.316 0.347
DBPedia 0.318 0.395 0.459
Robust04 0.407 0.406 0.492
Avg. 0.426 0.422 0.545
Table 6: Performance Comparison on BEIR Benchmarks (nDCG@10). The bold values highlight the best performance across the respective subsets.
Appendix G BRIGHT Evaluation
Datasets BM25 RankZephyr RaCT
Bio. 0.182 0.219 0.246
Earth. 0.279 0.237 0.331
Econ. 0.164 0.144 0.185
Psy. 0.134 0.103 0.228
Rob. 0.109 0.076 0.164
Stack. 0.163 0.137 0.157
Sus. 0.161 0.166 0.205
Pony 0.043 0.065 0.137
Leet. 0.247 0.247 0.253
AOPs 0.065 0.068 0.072
TheoT 0.021 0.020 0.082
TheoQ 0.073 0.073 0.081
Avg. 0.137 0.130 0.178
Table 7: Performance Comparison on BRIGHT Benchmarks (nDCG@10). The bold values highlight the best performance across the respective subsets.
Appendix H Examples of General LLM Tasks
To illustrate the degradation in general text understanding and generation capabilities of RankVicuna and Rankzephyr, we provide two examples below. Due to the length of RaCT’s response, we have omitted the middle portion of the answer.
H.1 Example 1
@@
RaCT Aloha and Mahalo: Discovering the Aloha Spirit in Hawaii As I stepped off the plane in Honolulu, I couldn’t help but feel a sense of excitement and anticipation wash over me. I had dreamed of visiting Hawaii for years, and finally, I was here. The warm tropical air enveloped me, and the sweet scent of plumeria filled my senses as I made my way to my hotel. Little did I know, this trip would be a journey not only to breathtaking natural wonders but also to a rich cultural heritage that would leave me feeling grateful and inspired. Immersing in the Local Culture: (Content omitted for brevity.) Must-See Attractions: (Content omitted for brevity.) A Taste of Aloha: (Content omitted for brevity.) Conclusion: (Content omitted for brevity.) Practical Tips and Recommendations: Visit Pearl Harbor and USS Arizona Memorial in the morning to avoid the crowds. Try local specialties like poke, laulau, and shave ice. Take a surf lesson or snorkeling trip to explore the islands’ stunning coastlines. Visit the Bishop Museum to learn about the history and geology of Hawaii. Explore the islands’ many farmers’ markets for fresh produce and local products. Learn traditional Hawaiian crafts, like leis and ukulele strings, from local artisans. Mahalo (thank you) for joining me on this journey to the Aloha State. I hope my story has inspired you to plan your own trip to Hawaii and experience the warmth and beauty of this incredible destination.
H.2 Example 2
@@
RaCT I’d be happy to help you write a resignation letter to your current employer. Here’s a sample template you can use as a guide: Dear [Recipient’s Name], I am writing to inform you of my decision to resign from my position as [Your Position] at [Company Name], effective [Date of Last Day of Work]. I want to express my deepest gratitude to you and the entire [Company Name] team for the opportunities I’ve had during my time here. (Content omitted for brevity.) Please don’t hesitate to reach out to me if you need anything after my departure. I can be reached at [Your Email Address] or [Your Phone Number]. Sincerely, [Your Signature] Some tips to keep in mind when writing your resignation letter: Be professional and respectful: Use a formal tone and language throughout the letter. Express gratitude: Let your employer know how much you appreciate the opportunities you’ve had and the experience you’ve gained. Offer to help with the transition: Show that you’re committed to making the transition as smooth as possible by offering to complete outstanding tasks and assist with the handover of your responsibilities. Keep it brief: Aim for a letter that is no more than one page in length. Proofread: Make sure to proofread your letter carefully to ensure it is free of errors and typos. Remember to customize the letter to fit your own circumstances and experiences. Good luck with your resignation, and I hope it leads to a positive and successful transition for you!